OpenStudy (anonymous):

Please help Simplify, rationalize all denominators: square root of 3x times square root of 5x

OpenStudy (anonymous):

\[\sqrt{3x} \times \sqrt{5x}\]

OpenStudy (phi):

When you multiply radicals you can put them together: \[\sqrt{3x}\sqrt{5x}= \sqrt{15x^{2}} \]

OpenStudy (phi):

Of course you can pull out the x^2 from the root sign

OpenStudy (anonymous):

thank you!